OpenStudy (anonymous):

find the soloutions of the system: y=x^2+3x-4 y=2x+2

OpenStudy (anonymous):

Start by rearranging the second equation: \[x = \frac{1}{2}(y-2)\] Substitute that into the first one and expand. What do you get?

OpenStudy (anonymous):

Or perhaps more simply, solve the equation: \[2x+2 = x^2+3x-4\]

OpenStudy (radar):

I like that MathBoy, Transpose getting \[x ^{2}+x-6\] can easily be factored and solved

OpenStudy (radar):

should of put "=0" at the end. lol

OpenStudy (radar):

clara_summer1, that you can do.

OpenStudy (radar):

Can you handle that now?

OpenStudy (radar):

@clara_summer1, are you sincerely interested in learning how to solve this?

OpenStudy (anonymous):

yes, i ve a child and it was very hard sitting down, sorry.

OpenStudy (radar):

O.K. Sorry, Do you know how to factor that in order to solve it? If not, I will show you.

OpenStudy (anonymous):

no, sorry

OpenStudy (radar):

\[x ^{2}+x-6=0\] (x+3)(x-2)=0 These are the factors, since they result in 0 one or both have to be 0 x+3=0 and x-2=0 x=-3 x=2

OpenStudy (anonymous):

ok so how to we get to the y?

OpenStudy (radar):

Factoring becomes easier with practice. but you are looking for factors that will result first in x^2 That is easy as it has to be xNow we have to find factors when mulltiplies will result in a -6, and when combined will result in a 1 as we have 1x.